3. Using the same amount of resources, USA and Canada can each produce oranges and salmonberries as shown in the following table, measured in thousands of bushels.

USA Canada
oranges salmonberries oranges salmonberries
12 0 6 0
3 3 3 3
0 4 0 6


a. Which country has the comparative advantage in producing salmonberries? Which one has the comparative advantage in producing oranges? Explain.

b. Suppose that each country is currently producing 3 thousand bushels of salmonberries and 3 thousand bushels of oranges. Show that both can be better off if they specialize in producing one good and then engage in trade.
1
Expert's answer
2019-05-31T09:06:19-0400

a) Canada has the comprative advantage in salmonberries, because its alternative cost of thousand bushels of salmonberry is 1 thousand bushels of orange (the country produces either 6 thousand bushels of oranges or 6 thousand bushels of salmonberries), while in the USA 1 thousand bushels of salmonberry=3 thousand bushels of oranges. The USA has the comprative advantage in oranges, because its alternative cost of thousand bushels of orange is 1/3 thousand bushels of salmonberry.

b) If Canada produces 6 thousand bushels of salmonberries, it can sell 3 thousand bushels of salmonberries in the USA and get 3*3=9 thousand bushels of oranges. Totally Canada will gain 9-3=6 thousand bushels of oranges.

If the USA produces 12 thousand bushels of oranges, it can sell 9 thousand bushels of oranges in Canada and get 9 thousand bushels of salmonberries. Totally the USA will gain 9-3=6 thousand bushels of salmonberries.